Why a Substance Can Be a Liquid at Room Temperature?

Calculate ΔS∘total for the evaporation of water at 25∘C.

H2O (l) ⇄ H2O (g)

ΔH∘f (kJ/mol) Sm∘ (J/mol*K)

H2O (l) - 285.830 69.91

H2O (g) - 241.818 188.83

ΔS∘ system?

ΔS∘surroundings?

ΔS∘total?
Answer #1

ΔS∘   = ΔS∘ products - ΔS∘ reactants

        = 188.83 - 69.91

         = 118.92 J / mol K

ΔS∘ system = 118.92 J / mol K

ΔH∘   = -241.818 - (-285.830)

         = 44.012 kJ / mol

ΔS∘surroundings   = - ΔH∘ / T

                               = -44012 / 298.15

                                = -147.62 J / mol . K

ΔS∘surroundings    = -147.62 J / mol . K

ΔS∘total   = ΔS∘ system + ΔS∘ surrounds

               = 118.92 - 147.62

              = -28.70 J / mol.K

ΔS∘total = - 28.70 J / mol.K

ΔS∘total   will be negative here . so that H2O will be existed as liquid at room temperature . it not is spontaneous to convert to gas at room temperature